OpenStudy (anonymous):

an alloy consist of three metal X,Y, Z calculate x:z x:y=2:3 and y:z=5:4

OpenStudy (anonymous):

x/y=2/3 implies that 3x=2y and y/z=5/4 implies that 4y=5z 2(2y)=5z now by substituting.....we get 2*3x=5z 6x=5z x/z=5/6